OpenStudy (anonymous):

Here's a question and the answer; how did they get there? Question: Find dm/dv for m=m_0/sqrt(1-(v^2/c^2) Answer: dm/dv= (m_0*v)/(c^2(1-(v^2/c^2))^3/2)

OpenStudy (kainui):

\[m=\frac{ m _{o} }{ \sqrt{1-\frac{ v^2 }{ c^2 }} }\] can be written as\[m= m _{o}(1-\frac{ v^2 }{ c^2 })^{-1/2}\]

OpenStudy (anonymous):

So does that mean they're assuming the \[v^2/c^2\] term is a constant?

OpenStudy (anonymous):

Since this isn't multivariable calculus, I doubt that we're intended to work with two constants. Perhaps their intention is that we treat that term AS the constant? Does that sound reasonable?

OpenStudy (kainui):

No, mo is constant, v is your variable. Morepresents the initial m value, and that's unchanging and constant, generally speaking.

OpenStudy (anonymous):

What do we do with the c^2 term then? Is that a constant also, simply by virtue of not being included as part of the dm/dv notation?

OpenStudy (kainui):

Yes, dm/dv means you're taking the derivative of m with respect to v. Similarly, when you take derivative of y with respect to x, that's just dy/dx. Just treat v like it was your "x"in a sense. Also,\[\frac{ v^2 }{ c^2 }=\frac{ 1 }{ c^2 }*v^2\] The reason I wrote it like that was to show how 1/c^2 is just like any other constant in front of your variable just like taking the derivative of 2x^2.

OpenStudy (anonymous):

That makes sense. I also just looked and realized I mistyped the answer slightly. Thank you very much for your help with that!

OpenStudy (kainui):

Cool, no problem.